Al a Mode, Inc., is considering one of two investment options. Option 1 is a $55,000 investment in new blending equipment that is expected to produce equal annual cash flows of $17,000 for each of seven years. Option 2 is a $62,000 investment in a new computer system that is expected to produce equal annual cash flows of $22,000 for each of five years. The residual value of the blending equipment at the end of the fifth year is estimated to be $11,000. The computer system has no expected residual value at the end of the fifth year.

the fifth year.

Present Value of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 0.890 0.826 0.797 0.756 0.694
3 0.840 0.751 0.712 0.658 0.579
4 0.792 0.683 0.636 0.572 0.482
5 0.747 0.621 0.567 0.497 0.402
6 0.705 0.564 0.507 0.432 0.335
7 0.665 0.513 0.452 0.376 0.279
8 0.627 0.467 0.404 0.327 0.233
9 0.592 0.424 0.361 0.284 0.194
10 0.558 0.386 0.322 0.247 0.162
Present Value of an Annuity of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 1.833 1.736 1.690 1.626 1.528
3 2.673 2.487 2.402 2.283 2.106
4 3.465 3.170 3.037 2.855 2.589
5 4.212 3.791 3.605 3.352 2.991
6 4.917 4.355 4.111 3.784 3.326
7 5.582 4.868 4.564 4.160 3.605
8 6.210 5.335 4.968 4.487 3.837
9 6.802 5.759 5.328 4.772 4.031
10 7.360 6.145 5.650 5.019 4.192

Assume there is sufficient capital to fund only one of the projects. Determine which project should be selected, comparing the (a) net present values and (b) present value indices of the two projects, assuming a minimum rate of return of 12%. Use the present value tables appearing above. a. Determine the net present values of the two projects. Blending Equipment Computer System Total present value of cash flows $ $ Less amount to be invested $ $ Net present value $ $ b. Determine the present value indices of the two projects. If required, round the present value index to two decimal places. Present Value Index Blending Equipment Computer System Which project should be selected? (If both present value indices are the same, either project will grade as correct.)
Answer #1
Computation of net present values of the two projects.
Blending Equipment
Total present value of cash flows = 17000*PVIFA(12%,5) + 11000*PVIF(12%,5)
Total present value of cash flows = 17000*3.605 + 11000*0.567
Total present value of cash flows = $67522
Less amount to be invested = $ 55000
Net present value = 67522 - 55000 = $12522
Computer System
Total present value of cash flows = 2200*PVIFA(12%,5)
Total present value of cash flows = 2200*3.605
Total present value of cash flows = $79310
Less amount to be invested = $ 62000
Net present value = 79310- 62000 = $17310
b. computation the present value indices of the two projects.
Blending Equipment
Present value index = Total present value of cash flows/Initial Investment
Present value index = 67522/55000
Present value index = 1.23
Computer System
Present value index = Total present value of cash flows/Initial Investment
Present value index =79310/62000
Present value index = 1.28
Project Computer System should be selected